Circular No. 9055

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF JUPITER A. A. Christou, Armagh Observatory, reports that his CCD photom-

etry of Jupiter V (Amalthea), obtained by F. Lewis (Faulkes Telescope Project, Open University) using the 2-m Faulkes Telescope South (+ Pan-STARRS z′ filter) reveals an eclipse of this satellite by Jupiter I (Io) on June 23 UT. Twenty-five 5-sec exposures were acquired between June 23.64799 and 23.66425. Amalthea was visible in all these frames with the exception of those starting at June 23.65317 and 23.65379 UT, consistent with a totality between June 23.65292 and 23.65401 predicted by utilizing a generic mutual-eclipse model (Christou 2005, Icarus 178, 171) in combination with the SPICE ephemeris kernel JUP230 for the Galilean satellites and Amalthea. Observations at high cadence (≪ 60 s) of further eclipses of Amalthea by Io are encouraged during the current Jovian equinox season. Predictions for July and August 2009 are given at website URL

http://www.arm.ac.uk/~aac/amalthea/.

COMETS C/2009 A7, C/2009 B8–B11, AND C/2009 C3–C4 (SOHO)

Further to IAUC 9042, additional apparent comets have been found on SOHO website images — all Kreutz sungrazers except for C/2009 B10 (Meyer group). C/2009 A7 was stellar in appearance (mag ∼ 7–7.5) in the C3 images; it was also imaged with the SECCHI HI-1A camera, which showed it to be small and slightly diffuse at mag ∼ 11. C/2009 B8 was slightly diffuse (mag ∼ 7). C/2009 B9 was stellar in appearance (mag ∼ 7.5). C/2009 B10 was stellar in appearance (mag ∼ 7.5) and perhaps slightly elongated. C/2009 B11 was also stellar in appearance (mag ∼ 7– 7.5). C/2009 C3 was bright (mag ∼ 2–3) with a thin tail in C3 images, while in C2 images it showed a long, thin tail that went past the field-of-view and showed no appreciable ‘head’. C/2009 C4 was slightly elongated, peaking at mag ∼ 7 (and brightening in the last image).

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 A7 Jan. 4.513 19^{h}14.^{m}0 −24^{o}40′ C3* AK 2009-M41
C/2009 B8  20.971 20 28.0 −21 46 C3  BZ 2009-M41
C/2009 B9  29.238 21 01.3 −20 01 C3  ZX 2009-M41
C/2009 B10  29.271 20 51.3 −16 28 C2  ZX 2009-M42
C/2009 B11  30.804 21 09.7 −18 51 C3  TH 2009-M42
C/2009 C3 Feb. 4.279 21 40.6 −18 41 C3/2 RM 2009-M42
C/2009 C4  7.936 21 52.8 −17 05 C3  BZ 2009-M54
2009 July 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green